# Byzantine Consensus in the Random Asynchronous Model

### v1

@@ 7. Related Work
Similar assumptions to the random asynchronous model have been explored by previous work. In message passing, Bracha and Toueg (Bracha and Toueg, 1985) define the fair scheduler, which ensures that in each message round, there is a non-zero constant probability that every correct process receives messages from the same set of correct processes. Under this scheduler, they proposed deterministic asynchronous binary consensus protocols for crash and Byzantine fault models. More recently, Tusk (Danezis et al., 2022) and Mahi-Mahi (Jovanovic et al., 2024) employ a form of random scheduling. Their random scheduler can be seen as a special case of ours: they only consider the $n=3f+1$ case and assume a standard round-based model with the subset of processes that a process “hears from” in a given round chosen uniformly at random among all possibilities. They leverage the random scheduler to increase the probability to commit at each round, and thus to reduce latency, whereas our paper focuses on circumventing impossibility results in standard asynchrony, at different ratios of fault-tolerance. They conduct experiments without randomization on a wide-area network, without observing loss of liveness, which can serve as motivation for our work.
@@
In shared memory, Aspnes (Aspnes, 2002) defines noisy scheduling, in which the adversary may chose the schedule, but that adversarial schedule is perturbed randomly. Under this assumption, deterministic asynchronous consensus becomes achievable. Also in shared memory, previous work introduce a stochastic scheduler (Alistarh et al., 2015, 2016), which schedules shared memory steps randomly. This line of work shows that many lock-free algorithms are essentially wait-free when run against a stochastic scheduler, because the schedules that would break wait-freedom have negligible probability.
Randomized consensus
@@
A large body of research leverages random coins to circumvent the FLP impossibility result (Fischer et al., 1983), which states that deterministic consensus is impossible in crash-prone asynchronous systems. In this approach, protocols relax deterministic termination to probabilistic termination, assuming processes have access a source of (cryptographically-secure) randomness that cannot be predicted by the adversary. Randomized consensus protocol employ either local coins (Ben-Or, 1983; Moniz et al., 2011; Zhang et al., 2023)—which produce randomness independently and locally at each process, without coordination with other processes—or common coins (Rabin, 1983; Cachin et al., 2000; Mostéfaoui et al., 2015; Danezis et al., 2022; Jovanovic et al., 2024; Spiegelman et al., 2022)—which, through the use of coordination and strong cryptographic primitives, ensure that all correct processes receive the same random output with some probability.
@@
Our algorithms for the $n=3f+1$ setting resemble existing coin-based random consensus protocols, with the randomness moved from the process logic to the schedule. In fact, all of the coin-based consensus protocols we examined can be transformed, with minor changes, into deterministic (coin-less) algorithms in the random asynchronous model. A natural question, then, is whether our model is equivalent to the standard asynchronous model with coin tosses. It is not: in the standard model, achieving safety whp when $n<3f$ is impossible, even if processes have access to randomness. This is because of a standard split-brain argument: the adversary can partition correct processes into two sets that never exchange messages, allowing Byzantine to force different decisions in each set. By contrast, our model makes long-lived network partitions occur with negligible probability, allowing safety whp even for $n<3f$.
Probabilistic quorum systems
@@
A line of work on probabilistic quorum systems (Malkhi et al., 2001; Yu, 2006) relaxes quorum intersection to be probabilistic rather than deterministic, and allows for probabilistic correctness guarantees. However, they are vulnerable to an adversarial scheduler (Aiyer et al., 2005). ProBFT (Avelas et al., 2024) addresses this through the use of verifiable random functions for quorum selection. These works are similar to ours in that correctness is probabilistic rather than deterministic, but their approach focuses on the $n=3f+1$ setting, and is mainly aimed at scalability and efficiency (e.g., communication complexity), whereas we aim to circumvent impossibilities in standard asynchrony across various fault tolerance ratios.
8. Conclusion
@@
We introduce the random asynchronous model, a novel relaxation of the classic asynchronous model that replaces adversarial message scheduling with a randomized scheduler. By eliminating the adversary’s ability to indefinitely delay messages, our model circumvents traditional impossibility results in asynchronous Byzantine consensus while preserving unbounded message delays and tolerating Byzantine faults. Our approach avoids the need for synchronized periods (as in partial synchrony) or cryptographic randomness (as in randomized consensus), offering a foundation for practical alternatives to existing asynchronous systems. We demonstrated that this relaxation enables new feasibility results across different resilience thresholds: deterministic safety and probabilistic termination for $n=3f+1$, deterministic termination with safety holding with high probability (whp) for $n=2f+1$, and weak validity with whp agreement for $n=f+2$. These results are complemented by impossibility bounds, showing our protocols achieve near-optimal guarantees under the model.
@@
Future work could explore extensions of this model to other distributed computing problems, such as state machine replication, and investigate empirical performance trade-offs in real-world deployments. By bridging the gap between theoretical impossibility and practical assumptions, we believe our model opens avenues for efficient, resilient consensus protocols.
@@ A.1. Modeling Challenge
  1. (1)
    A round-based model, similar to the fair scheduler model of Bracha and Toueg (Bracha and Toueg, 1985). In each communication round, a correct process sends a message to every process and waits to hear back from $n-f$ processes. The random scheduler assumption is: in each round, a correct process has a non-trivial (i.e., lower-bounded by a constant) probability of hearing from any subset of $n-f$ processes. This model has the advantage of being easy to work with, but is too restrictive, as it restricts algorithms to the round-based structure.
  2. (2)
@@ A.2. Algorithmic Challenge
  • Thus, no correct processes is aware of 00 as a valid input value before round $R$.
@@
This algorithm illustrates the main challenge of designing correct algorithms under the random asynchronous model when $n\leq 2f+1$: even if the model ensures that all correct processes communicate with each other eventually, Byzantine processes can still equivocate and correct processes do not know necessarily know which messages are from correct processes and which are not.
@@ Appendix B Proofs for Deterministic Safety Algorithms
In our proofs we rely on the following properties of Byzantine Reliable Broadcast (BRB) (Cachin et al., 2011):
BRB-Validity:
@@
BRB-Consistency:
If some correct process delivers a message $m$ and another correct process delivers a message $m$, then $m=m$.
BRB-Totality:
@@
Proof.
If all correct processes propose the same value $v$, then at least $12f+$ 1 processes BRB-broadcast an Init message for $v$, and therefore at most $f$ processes BRB-broadcast an Init message for $1-v$. Thus $v$ will be the majority value among all Init messages delivered in phase 1, at all correct processes. Thus all correct processes will BRB-broadcast an Echo message for $v$. Furthermore, no Byzantine process can produce a valid Echo message for $1-v$, since to do so would require a set of $12f+$ 1 Init message with a majority value of $1-v$. This is impossible due to the properties of BRB and the fact that at most $f$ processes have BRB-broadcast an Init message for $1-v$. So, all valid Echo messages received by correct processes will be for $v$, so all correct processes will commit $v$ at line 10. ∎
@@
Proof.
If a correct process $p_{1}$ commits $v$ at line 10, then it must have delivered a set $S_{1}$ of $12f+$ 1 Echo messages for $v$ at line 8. Take now another process $p_{2}$ and consider the set $S_{2}$ of $12f+$ 1 Echo messages it delivers at line 8. By quorum intersection, $S_{1}$ and $S_{2}$ must intersect in at least $f+1$ messages. By the BRB-Consistency property, these $f+1$ messages must be identical at $p_{1}$ and $p_{2}$. Thus $p_{2}$ delivers at least $f+1$ Echo messages for $v$, which constitutes a majority of the $12f+$ 1 Echo messages it delivers overall. So if $p_{2}$ commits a value at line 10, then it must commit $v$, and if $p_{2}$ adopts a value at line 13, then it must adopt $v$. ∎
@@
Proof.
Follows immediately from the algorithm and from the properties of Byzantine Reliable Broadcast. Processes perform two phases; the only blocking step of each phase is waiting for $n-f$ messages (lines 3 and 8). This waiting eventually terminates, by the BRB-Validity property and the fact that there are at least $n-f$ correct processes. ∎
@@
Proof.
This follows from the strong validity property of the Round procedure (Lemma B.1): if all correct processes propose $v$ to consensus, then all correct processes propose $v$ to Round in the first round, where by Lemma B.1, all correct processes commit $v$, and thus all correct processes decide $v$ at line 6. ∎
@@
For $p$ to decide $v$ at round $r$, Round must output $(\textsc{Commit},v)$ in that round. Thus, by the consistency property of Round, $\textsc{Round}(r,\cdot)$ must output $(\cdot,v)$ at all correct processes. If $\textsc{Round}(r,\cdot)$ outputs $(\textsc{Commit},v)$ for $p^{\prime}$, then $p^{\prime}$ decides $v$ at round $r$ (line 6). Otherwise, all correct processes input $v$ to $\textsc{Round}(r+1,\cdot)$, and by the strong validity property, all processes (including $p^{\prime}$) will output $(\textsc{Commit},v)$ and decide $v$ at round $r+1$. ∎
@@
Proof.
We can describe the execution of the protocol as a Markov chain with states $10,\ldots,n-f=2f+$ 1; the system is at state $i$ if $i$ correct processes have estimate ($est_{i}$ variable) equal to 00 before invoking Round. Due to the strong validity property of the Round procedure, states 00 and $12f+$ 1 are absorbing states. There is a non-zero transition probability from each state (including 00 and $12f+$ 1), to state 00 or $12f+$ 1, or both (we show this below). Therefore, with probability $1$, the system will eventually reach one of the two absorbing states and remain there. Once this happens (i.e., once all processes have the same $est_{i}$ variable), the strong validity property of Round ensures that all processes (who have not decided yet) will decide within a round.
@@
  • $i<f+1$: in this case 00 is the minority value among correct processes. In schedule $S$, the $n-f$ Init messages delivered by correct process at line 3 are all from correct processes. Thus, every correct process sees $i$ 00s and $2f+1-i$ $1$; $1$ is the majority value, so all correct processes adopt it for phase 2. In phase 2, $S$ again ensures that the $n-f$ Echo messages delivered by correct process at line 8 are all from correct processes. Thus, all correct processes see $12f+$ 1 Echo messages for $1$ and commit $1$, bringing the system to state 00.
@@ C.2. Algorithm
The purpose of the second phase is to verify if $v^{*}$ was also observed by enough other processes. After a process receives $n-f$ phase-two messages (line 8), it checks if more than $f$ phase-two proposals are equal to $v^{*}$, and if so commits $v^{*}$ (line 11). A process adopts $v^{*}$ if it receives $v^{*}$ in the second phase, but is unable to collect enough $v^{*}$ values to decide (line 13). Finally, it is possible that a process does not receive $v^{*}$ in the second phase (either because no such value was found in phase one or simply because it has received only $\bot$ in phase two); in this case the process adopts the fist value it received in phase one (line 15).
As in the Byzantine case, the main intuition is that at each round, a favorable schedule can help processes agree (i.e., adopt the same estimate), by causing them to adopt the same estimate at line 15 (e.g., by ensuring that the first Init message they receive is from the same process). Since the schedule is random, it has a non-zero chance of being favorable at each round. Thus, with probability $1$, the schedule will eventually be favorable.
C.3. Proofs
@@
Proof.
We say that a value $v$ is valid if it is the input value of some process. We want to show that processes only return valid values. We observe that (1) Init messages only contain valid values (line 2), and therefore (2) Echo messages only contain valid values or $\bot$ (lines 48). If a process $p$ returns $v$ at line 11 or 13, then $p$ received at least one Echo message for $v$, and thus $v$ is valid by (2) above. If $p$ returns $v$ at line 15, then $p$ received at least one Init message for $v$, and thus $v$ is valid by (1) above. ∎
@@
Proof.
If all processes propose the same value $v$, then all processes send $\langle\textsc{Init},v\rangle$ at line 2; all processes receive at least $n/2$ $\langle\textsc{Init},v\rangle$ messages (since $n-f=f+1\geq n/2$); all processes adopt $v$ as their proposal for the second phase and send $\langle\textsc{Echo},v\rangle$ at line 8; all processes receive $n-f=f+1$ $\langle\textsc{Echo},v\rangle$ and return $\langle\textsc{Commit},v\rangle$. ∎
@@
Proof.
Assume the lemma does not hold. Then $p$ must have received more than $n/2$ $\langle\textsc{Init},v\rangle$ messages, and some process $p^{\prime}$ must have received more than $n/2$ $\langle\textsc{Init},v^{\prime}\rangle$ for some $v^{\prime}\neq v$. By quorum intersection, it follows that some process must have sent both an $\langle\textsc{Init},v\rangle$ and $\langle\textsc{Init},v^{\prime}\rangle$ message. This is a contradiction, as processes only send one Init message at line 2. ∎
@@
Proof.
If process $p$ returns $\langle\textsc{Commit},v\rangle$, it must do so at line 11, after having received $f+1$ $\langle\textsc{Echo},v\rangle$ messages. Therefore, every other process $p^{\prime}$ that returns, must receive at least one $\langle\textsc{Echo},v\rangle$ message. If $p^{\prime}$ also receives $f+1$ $\langle\textsc{Echo},v\rangle$ messages, then it returns $\langle\textsc{Commit},v\rangle$. Otherwise, by Lemma C.3, $p^{\prime}$ cannot receive an Echo message for any other value $v^{\prime}\neq v$, so $p^{\prime}$ returns $\langle\textsc{Adopt},v\rangle$ at line 13. ∎
@@
Proof.
This follows immediately from the construction of the algorithm. Processes perform two phases; the only blocking step of each phase is waiting for $n-f$ messages (lines 3 and 9). This waiting eventually terminates, since there are at least $n-f$ correct processes. ∎
@@
Proof.
We prove the result by induction on the round $r$. For the base case: If $r=0$, then all processes input their consensus inputs into Round. For the induction case: assume the lemma holds up to $r>0$, and consider the case of $r+1$. By the induction hypothesis and the integrity property of the Round procedure, any output of $\textsc{Round}(r,\cdot)$ must be the consensus input of some process. Since the input of $\textsc{Round}(r+1,\cdot)$ is the output of $\textsc{Round}(r,\cdot)$, the lemma must also hold for the case of $r+1$. This completes the induction. ∎
Proof of Theorem C.6.
If a process $p$ decides a value $v$, then the Round procedure must have output $(\textsc{Commit},v)$ at line 6. By Lemma C.7 and the integrity property of the Round procedure, it follows that $v$ is the consensus input of some process. ∎
@@
For $p$ to decide $v$ at round $r$, Round must output $(\textsc{Commit},v)$ in that round. Thus, by the consistency property of Round, $\textsc{Round}(r,\cdot)$ must output $(\cdot,v)$ at all processes. If $\textsc{Round}(r,\cdot)$ outputs $(\textsc{Commit},v)$ for $p^{\prime}$, then $p^{\prime}$ decides $v$ at round $r$ (line 6). Otherwise, no other value than $v$ can be input to $\textsc{Round}(r+1,\cdot)$, and by the strong validity property, all processes (including $p^{\prime}$) will output $(\textsc{Commit},v)$ and decide $v$ at round $r+1$. ∎
@@
With crash faults, Algorithm 1 satisfies termination with probability $1$.
Proof.
We can describe the execution of the protocol as a Markov chain with states $10,\ldots,2f+$ 1; the system is at state $i$ if $i$ processes have estimate ($est_{i}$ variable) equal to 00 before invoking Round. Due to the strong validity property of the Round procedure, states 00 and $12f+$ 1 are absorbing states. There is a non-zero transition probability from each state, other than 00 and $12f+$ 1, to each other state (we show this below). Therefore, with probability $1$, the system will eventually reach one of the two absorbing states and remain there. Once this happens (all processes have the same $est_{i}$ variable), the strong validity property of Round ensures that all processes (who have not decided yet) will decide within a round.

### v2

@@ 8 Related Work
Random scheduling. Similar assumptions to the random asynchronous model have been explored by previous work. In message passing, Bracha and Toueg [8] define the fair scheduler, which ensures that in each message round, there is a non-zero constant probability that every correct process receives messages from the same set of correct processes. Under this scheduler, they proposed deterministic asynchronous binary consensus protocols for crash and Byzantine fault models. More recently, Tusk [11] and Mahi-Mahi [16] employ a form of random scheduling. Their random scheduler can be seen as a special case of ours: they only consider the $n=3f+1$ case and assume a standard round-based model with the subset of processes that a process “hears from” in a given round chosen uniformly at random among all possibilities. They leverage the random scheduler to increase the probability to commit at each round, and thus to reduce latency, whereas our paper focuses on circumventing impossibility results in standard asynchrony, at different ratios of fault-tolerance. They conduct experiments without randomization on a wide-area network, without observing loss of liveness, which can serve as motivation for our work.
@@
In shared memory, Aspnes [4] defines noisy scheduling, in which the adversary may choose the schedule, but that adversarial schedule is perturbed randomly. Under this assumption, deterministic asynchronous consensus becomes achievable. Also in shared memory, previous work introduce a stochastic scheduler [3, 2], which schedules shared memory steps randomly. This line of work shows that many lock-free algorithms are essentially wait-free when run against a stochastic scheduler, because the schedules that would break wait-freedom have negligible probability.
@@
Randomized consensus. A large body of research leverages random coins to circumvent the FLP impossibility result [13], which states that deterministic consensus is impossible in crash-prone asynchronous systems. In this approach, protocols relax deterministic termination to probabilistic termination, assuming processes have access to a source of (cryptographically-secure) randomness that cannot be predicted by the adversary. Randomized consensus protocols employ either local coins [7, 21, 28]—which produce randomness independently and locally at each process, without coordination with other processes—or common coins [23, 10, 22, 11, 16, 25]—which, through the use of coordination and strong cryptographic primitives, ensure that all correct processes receive the same random output with some probability.
@@
Our algorithms for the $n=3f+1$ setting resemble existing coin-based random consensus protocols, with the randomness moved from the process logic to the schedule. In fact, all of the coin-based consensus protocols we examined can be transformed, with minor changes, into deterministic (coin-less) algorithms in the random asynchronous model. A natural question, then, is whether our model is equivalent to the standard asynchronous model with coin tosses. It is not: in the standard model, achieving safety w.h.p. when $n<3f$ is impossible, even if processes have access to randomness. This is because of a standard network partition argument: the adversary can partition correct processes into two sets that never exchange messages, allowing Byzantine to force different decisions in each set. By contrast, our model makes long-lived network partitions occur with negligible probability, allowing safety w.h.p. even for $n<3f$.
@@
Probabilistic quorum systems. A line of work on probabilistic quorum systems [20, 27] relaxes quorum intersection to be probabilistic rather than deterministic, and allows for probabilistic correctness guarantees. However, they are vulnerable to an adversarial scheduler [1]. ProBFT [5] addresses this through the use of verifiable random functions for quorum selection. These works are similar to ours in that correctness is probabilistic rather than deterministic, but their approach focuses on the $n=3f+1$ setting, and is mainly aimed at scalability and efficiency (e.g., communication complexity), whereas we aim to circumvent impossibilities in standard asynchrony across various fault tolerance ratios.
9 Conclusion
@@
We introduce the random asynchronous model, a novel relaxation of the classic asynchronous model that replaces adversarial message scheduling with a randomized scheduler. By eliminating the adversary’s ability to indefinitely delay messages, our model circumvents traditional impossibility results in asynchronous Byzantine consensus while preserving unbounded message delays and tolerating Byzantine faults. Our approach avoids the need for synchronized periods (as in partial synchrony) or cryptographic randomness (as in randomized consensus), offering a foundation for practical alternatives to existing asynchronous systems. We demonstrated that this relaxation enables new feasibility results across different resilience thresholds: deterministic safety and probabilistic termination for $n=3f+1$, deterministic termination with safety holding with high probability (w.h.p.) for $n=2f+1$, and weak validity with w.h.p. agreement for $n=f+2$. These results are complemented by impossibility bounds, showing our protocols achieve near-optimal guarantees under the model.
@@
Future work could explore extensions of this model to other distributed computing problems, such as state machine replication, and investigate empirical performance trade-offs in real-world deployments. By bridging the gap between theoretical impossibility and practical assumptions, we believe our model opens avenues for more efficient and resilient distributed protocols.
@@ A.1 Modeling Challenge
  1. 1.
    A round-based model, similar to the fair scheduler model of Bracha and Toueg [8]. In each communication round, a correct process sends a message to every process and waits to hear back from $n-f$ processes. The random scheduler assumption is: in each round, a correct process has a non-trivial (i.e., lower-bounded by a constant) probability of hearing from any subset of $n-f$ processes. This model has the advantage of being easy to work with, but is too restrictive, as it restricts algorithms to the round-based structure.
  2. 2.
@@ A.2 Algorithmic Challenge
  • Thus, no correct process is aware of 00 as a valid input value before round $R$.
@@
This algorithm illustrates the main challenge of designing correct algorithms under the random asynchronous model when $n\leq 2f+1$: even if the model ensures that all correct processes communicate with each other eventually, Byzantine processes can still equivocate and correct processes do not necessarily know which messages are from correct processes and which are not.
@@ Appendix B Proofs for Deterministic Safety Algorithms
In our proofs we rely on the following properties of Byzantine Reliable Broadcast (BRB) [9]:
BRB-Validity
@@
BRB-Consistency
If some correct process delivers a message $m$ and another correct process delivers a message $m^{\prime}$, then $m=m^{\prime}$.
BRB-Totality
@@
Proof B.2.
If all correct processes propose the same value $v$, then at least $12f+$ 1 processes BRB-broadcast an Init message for $v$, and therefore at most $f$ processes BRB-broadcast an Init message for $1-v$. Thus $v$ will be the majority value among all Init messages delivered in phase 1, at all correct processes. Thus all correct processes will BRB-broadcast an Echo message for $v$. Furthermore, no Byzantine process can produce a valid Echo message for $1-v$, since to do so would require a set of $12f+$ 1 Init message with a majority value of $1-v$. This is impossible due to the properties of BRB and the fact that at most $f$ processes have BRB-broadcast an Init message for $1-v$. So, all valid Echo messages received by correct processes will be for $v$, so all correct processes will commit $v$ at line 9.
@@
Proof B.4.
If a correct process $p_{1}$ commits $v$ at line 9, then it must have delivered a set $S_{1}$ of $12f+$ 1 Echo messages for $v$ at line 7. Take now another process $p_{2}$ and consider the set $S_{2}$ of $12f+$ 1 Echo messages it delivers at line 7. By quorum intersection, $S_{1}$ and $S_{2}$ must intersect in at least $f+1$ messages. By the BRB-Consistency property, these $f+1$ messages must be identical at $p_{1}$ and $p_{2}$. Thus $p_{2}$ delivers at least $f+1$ Echo messages for $v$, which constitutes a majority of the $12f+$ 1 Echo messages it delivers overall. So if $p_{2}$ commits a value at line 9, then it must commit $v$, and if $p_{2}$ adopts a value at line 12, then it must adopt $v$.
@@
Proof B.6.
Follows immediately from the algorithm and from the properties of Byzantine Reliable Broadcast. Processes perform two phases; the only blocking step of each phase is waiting for $n-f$ messages (lines 3 and 7). This waiting eventually terminates, by the BRB-Validity property and the fact that there are at least $n-f$ correct processes.
@@
Proof B.8.
This follows from the strong validity property of the Round procedure (Lemma B.1): if all correct processes propose $v$ to consensus, then all correct processes propose $v$ to Round in the first round, where by Lemma B.1, all correct processes commit $v$, and thus all correct processes decide $v$ at line 6.
@@
For $p$ to decide $v$ at round $r$, Round must output $(\textsc{Commit},v)$ in that round. Thus, by the consistency property of Round, $\textsc{Round}(r,\cdot)$ must output $(\cdot,v)$ at all correct processes. If $\textsc{Round}(r,\cdot)$ outputs $(\textsc{Commit},v)$ for $p^{\prime}$, then $p^{\prime}$ decides $v$ at round $r$ (line 6). Otherwise, all correct processes input $v$ to $\textsc{Round}(r+1,\cdot)$, and by the strong validity property, all processes (including $p^{\prime}$) will output $(\textsc{Commit},v)$ and decide $v$ at round $r+1$.
@@
Proof B.12.
We can describe the execution of the protocol as a Markov chain with states $10,\ldots,n-f=2f+$ 1; the system is at state $i$ if $i$ correct processes have estimate ($est_{i}$ variable) equal to 00 before invoking Round. Due to the strong validity property of the Round procedure, states 00 and $12f+$ 1 are absorbing states. There is a non-zero transition probability from each state (including 00 and $12f+$ 1), to state 00 or $12f+$ 1, or both (we show this below). Therefore, almost surely, the system will eventually reach one of the two absorbing states and remain there. Once this happens (i.e., once all processes have the same $est_{i}$ variable), the strong validity property of Round ensures that all processes (who have not decided yet) will decide within a round.
@@
  • $i<f+1$: in this case 00 is the minority value among correct processes. In schedule $S$, the $n-f$ Init messages delivered by correct processes at line 3 are all from correct processes. Thus, every correct process sees $i$ 00s and $2f+1-i$ $1$s; $1$ is the majority value, so all correct processes adopt it for phase 2. In phase 2, $S$ again ensures that the $n-f$ Echo messages delivered by correct processes at line 7 are all from correct processes. Thus, all correct processes see $12f+$ 1 Echo messages for $1$ and commit $1$, bringing the system to state 00.
@@ C.2 Algorithm
The purpose of the second phase is to verify if $v^{*}$ was also observed by enough other processes. After a process receives $n-f$ phase-two messages (line 7), it checks if more than $f$ phase-two proposals are equal to $v^{*}$, and if so commits $v^{*}$ (line 11). A process adopts $v^{*}$ if it receives $v^{*}$ in the second phase, but is unable to collect enough $v^{*}$ values to decide (line 13). Finally, it is possible that a process does not receive $v^{*}$ in the second phase (either because no such value was found in phase one or simply because it has received only $\bot$ in phase two); in this case the process adopts the fist value it received in phase one (line 15).
As in the Byzantine case, the main intuition is that at each round, a favorable schedule can help processes agree (i.e., adopt the same estimate), by causing them to adopt the same estimate at line 15 (e.g., by ensuring that the first Init message they receive is from the same process). Since the schedule is random, it has a non-zero chance of being favorable at each round. Thus, almost surely, the schedule will eventually be favorable.
C.3 Proofs
@@
Proof C.2.
We say that a value $v$ is valid if it is the input value of some process. We want to show that processes only return valid values. We observe that (1) Init messages only contain valid values (line 2), and therefore (2) Echo messages only contain valid values or $\bot$ (lines 48). If a process $p$ returns $v$ at line 11 or 13, then $p$ received at least one Echo message for $v$, and thus $v$ is valid by (2) above. If $p$ returns $v$ at line 15, then $p$ received at least one Init message for $v$, and thus $v$ is valid by (1) above.
@@
Proof C.4.
If all processes propose the same value $v$, then all processes send $\langle\textsc{Init},v\rangle$ at line 2; all processes receive at least $n/2$ $\langle\textsc{Init},v\rangle$ messages (since $n-f=f+1\geq n/2$); all processes adopt $v$ as their proposal for the second phase and send $\langle\textsc{Echo},v\rangle$ at line 8; all processes receive $n-f=f+1$ $\langle\textsc{Echo},v\rangle$ and return $\langle\textsc{Commit},v\rangle$.
@@
Proof C.6.
Assume the lemma does not hold. Then $p$ must have received more than $n/2$ $\langle\textsc{Init},v\rangle$ messages, and some process $p^{\prime}$ must have received more than $n/2$ $\langle\textsc{Init},v^{\prime}\rangle$ for some $v^{\prime}\neq v$. By quorum intersection, it follows that some process must have sent both an $\langle\textsc{Init},v\rangle$ and $\langle\textsc{Init},v^{\prime}\rangle$ message. This is a contradiction, as processes only send one Init message at line 2.
@@
Proof C.8.
If process $p$ returns $\langle\textsc{Commit},v\rangle$, it must do so at line 11, after having received $f+1$ $\langle\textsc{Echo},v\rangle$ messages. Therefore, every other process $p^{\prime}$ that returns, must receive at least one $\langle\textsc{Echo},v\rangle$ message. If $p^{\prime}$ also receives $f+1$ $\langle\textsc{Echo},v\rangle$ messages, then it returns $\langle\textsc{Commit},v\rangle$. Otherwise, by Lemma C.5, $p^{\prime}$ cannot receive an Echo message for any other value $v^{\prime}\neq v$, so $p^{\prime}$ returns $\langle\textsc{Adopt},v\rangle$ at line 13.
@@
Proof C.10.
This follows immediately from the construction of the algorithm. Processes perform two phases; the only blocking step of each phase is waiting for $n-f$ messages (lines 3 and 9). This waiting eventually terminates, since there are at least $n-f$ correct processes.
@@
Proof C.13.
We prove the result by induction on the round $r$. For the base case: If $r=0$, then all processes input their consensus inputs into Round. For the induction case: assume the lemma holds up to $r>0$, and consider the case of $r+1$. By the induction hypothesis and the integrity property of the Round procedure, any output of $\textsc{Round}(r,\cdot)$ must be the consensus input of some process. Since the input of $\textsc{Round}(r+1,\cdot)$ is the output of $\textsc{Round}(r,\cdot)$, the lemma must also hold for the case of $r+1$. This completes the induction.
Proof C.14 (Proof of Theorem C.11).
If a process $p$ decides a value $v$, then the Round procedure must have output $(\textsc{Commit},v)$ at line 6. By Lemma C.12 and the integrity property of the Round procedure, it follows that $v$ is the consensus input of some process.
@@
For $p$ to decide $v$ at round $r$, Round must output $(\textsc{Commit},v)$ in that round. Thus, by the consistency property of Round, $\textsc{Round}(r,\cdot)$ must output $(\cdot,v)$ at all processes. If $\textsc{Round}(r,\cdot)$ outputs $(\textsc{Commit},v)$ for $p^{\prime}$, then $p^{\prime}$ decides $v$ at round $r$ (line 6). Otherwise, no other value than $v$ can be input to $\textsc{Round}(r+1,\cdot)$, and by the strong validity property, all processes (including $p^{\prime}$) will output $(\textsc{Commit},v)$ and decide $v$ at round $r+1$.
@@
With crash faults, Algorithm 1 satisfies termination almost surely.
Proof C.18.
We can describe the execution of the protocol as a Markov chain with states $10,\ldots,2f+$ 1; the system is at state $i$ if $i$ processes have estimate ($est_{i}$ variable) equal to 00 before invoking Round. Due to the strong validity property of the Round procedure, states 00 and $12f+$ 1 are absorbing states. There is a non-zero transition probability from each state, other than 00 and $12f+$ 1, to each other state (we show this below). Therefore, almost surely, the system will eventually reach one of the two absorbing states and remain there. Once this happens (all processes have the same $est_{i}$ variable), the strong validity property of Round ensures that all processes (who have not decided yet) will decide within a round.